OpenStudy (anonymous):

Let f(x) =e^(x/2) . If the second degree Taylor polynomial for f about x = 0 is used to approximate f on the interval [0, 2], what is the Lagrange error bound for the maximum error on the interval [0, 2]?

OpenStudy (anonymous):

\[e ^{x/2}\]

OpenStudy (anonymous):

By Definition the error is \[\Large \frac{x^3 f^{(3)}(c)}{3!}=\frac{1}{48 } e^{c/2} x^3 \] What is its maximum on the interval [0,2]

OpenStudy (anonymous):

The maximum occurs when c=x=2\[ \Large \frac{1}{48 } e^{2/2} 2^3\approx 0.453047 \]

OpenStudy (anonymous):

I'm confused, 1/48?

OpenStudy (anonymous):

\[ \Large f^{(3)}(x)=\frac{e^{x/2}}{8}\\ 8 (3!)=8(6)=48 \]

OpenStudy (anonymous):

oh, I see. Thank you very much :]

OpenStudy (anonymous):

YW